OpenStudy (anonymous):

Solve , where is a real number. Simplify your answer as much as possible. (y+5)exponent2-18=0 help please. If there is more than one solution, separate them with commas. If there is no solution, click "No solution."

OpenStudy (anonymous):

is the equation (y=5)^2-18=0?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no its (y+5)^2-18=0

OpenStudy (anonymous):

ok. (y+5)^2-18=0 expand (y+5)^2

OpenStudy (anonymous):

Y^2+10x+25-18=0

OpenStudy (anonymous):

y^2+10x+7

OpenStudy (anonymous):